OpenStudy (anonymous):

solve using ratios: the ratio of boys to girls in a sophomore class of 460 students is 11:12. Find the number of boys. I haven't done ratios since 6th grade O.o please help! Thank you:)

OpenStudy (kropot72):

If there are n items of two kinds, type a and type b, the number of type a items is found from: \[n \times \frac{a}{a+b}\] Can you do the problem now?

OpenStudy (anonymous):

253 boys?

OpenStudy (anonymous):

or is it 220 boys? .....

hero (hero):

Hint: Add 11 + 12 to get the total number of boys and girls then find x of this proportion: \(\large\frac{11}{23} = \frac{x}{460}\)

OpenStudy (anonymous):

thank you so it is 220 :)